OpenStudy (anonymous):

2(3k+4)-(9k-7)=6 what does k equal?

OpenStudy (precal):

distribute the 2 and the -1

OpenStudy (anonymous):

how did u get the -1

OpenStudy (precal):

no distribute the -1

OpenStudy (anonymous):

2*3k+4*2-9k+7=6 6k+8-9k+7=6 -3k+15=6 -3k=-9 k=3

OpenStudy (anonymous):

where sis u get the -1 from

OpenStudy (precal):